OpenStudy (anonymous):

stuck on this problem: 0.5b + 4 = 2(b + 2) :/

OpenStudy (anonymous):

0.5b + 4 = 2(b+2) 0.5b + 4 = 2b + 4 0.5b = 2b 0 = 1.5b b = 0

OpenStudy (anonymous):

thank you! :)